OpenStudy (anonymous):

Verify the identity. cot(x-pi/2) =-tanx

OpenStudy (anonymous):

i don't understand how to work out cot(x-pi/2)

OpenStudy (anonymous):

cot(x-pi/2)= cos(x-pi/2)/sin(x-p/2) ; (1/ what the tan would have been) cos(x-pi/2)=cosxcos (pi/2)+sinxsin(pi/2) = cosx(0)+sin(1) = sinx if we simplify sin(x-pi/2) the same way, we are left with -cosx So the first step simplifies to -sinx/cosx Which is another way to write -(tanx)

OpenStudy (anonymous):

thankyou!